# A Universal Deep-Learning Framework for Materials X-ray Absorption Spectra

### v2

@@ Abstract
X-ray absorption spectroscopy (XAS) is a powerful characterization technique for probing the local chemical environment of absorbing atoms. However, analyzing XAS data presents with significant challenges, often requiring extensive, computationally intensive simulations, as well as significant domain expertise. These limitations hinder the development of fast, robust XAS analysis pipelines that are essential in high-throughput studies and for autonomous experimentation. We address these challenges with OmniXAS, a framework that contains a suite of transfer learning approaches for XAS prediction, each uniquely contributing to improved accuracy and efficiency, as demonstrated on K-edge spectra database covering eight 3d transition metals (Ti – Cu). The OmniXAS framework is built upon three distinct strategies. First, we use M3GNet [Nat. Comput. Sci., 2, 718 (2022)] to derive latent representations of the local chemical environment of absorption sites as input for XAS prediction, achieving up to order-of-magnitude improvements over conventional featurization techniques. Second, we employ a hierarchical transfer learning strategy, training a universal multi-task model across elements before fine-tuning for element-specific predictions. Models based on this cascaded approach after element-wise fine-turning outperform element-specific models by up to 69%. Third, we implement cross-fidelity transfer learning, adapting a universal model to predict spectra generated by simulation of a different fidelity with a much higher computational cost. This approach improves prediction accuracy by up to 11% over models trained on the target fidelity alone. Our approach significantly boosts the throughput of XAS modeling by orders of magnitude as compared to first-principles simulations and is extendable to XAS prediction for a broader range of elements. The proposed transfer learning framework is generalizable to enhance deep-learning models that target other properties in materials research.
@@ I Introduction
In this work, we address the need for a generalizable ML framework in the XANES domain. Our approach leverages transfer learning and domain adaptation techniques to develop universal models. By adopting M3GNet-derived latent features [28] to predict K-edge XANES spectra for a range of 3d transition metals, we demonstrate that this method not only outperforms traditional featurization techniques but also effectively bridges elements and fidelities, providing a scalable solution for XANES analysis across a wide range of materials. The OmniXAS framework lays the groundwork for future applications in high-throughput and real-time XAS analysis, significantly reducing the reliance on computationally expensive simulations.
@@ II.1 Data Acquisition and Curation
Before ML models can be trained, structure and spectral data need to be sourced, curated, and standardized. We used the Lightshow Python package [13, 10] to pull all available Materials Project structural data of pure compounds, and oxides containing at most 0, 1 or 2 other unique atom types via Pymatgen [31] for Ti, V, Cr, Mn, Fe, Co, Ni and Cu. FEFF9 and VASP spectral input files were also written using Lightshow.
@@ II.3 Transfer Learning
Based on the categorization proposed by Pan and Yang [46], transfer learning is classified according to the approaches used, the level of similarity between source and target tasks, and the similarity of input (domain) data. In this paper, we focus on two of the categories they define: inductive-transfer-learning and domain-adaptation. For these categories, we employ the feature-transfer and fine-tuning approaches, respectively.
@@ II.4.1 M3GNet block
M3GNet is a graph deep learning model designed to predict the inter-atomic potential, forces, and stresses from the material graph representation [28]. It was trained on a comprehensive dataset from the Materials Project, comprising over 17 million entries and encompassing 89 elements from the periodic table. The efficiency M3GNet was demonstrated by the study that predicted almost 2 million stable materials from a total of approximately $32$ million candidate structures [28].
@@ II.4.3 XAS block
Here, $\hat{\bm{\mu}}\in\mathbb{R}^{141}$ represents the predicted XAS spectrum across 141 energy grid points.
Each component of the MLP serves a specific purpose in the learning process, contributing to the model’s overall effectiveness:
  • Linear layers form the backbone of the MLP, enabling the network to learn complex mappings between input and output spaces:
    $\mathrm{Linear}_{l}(\mathbf{h})=W_{l}\mathbf{h}+\mathbf{b}_{l}.$ (7)
    Here, the weight matrix, $W_{l},$ and bias vector, $\mathbf{b}_{l},$ are learnable parameters that the network adjusts during training to capture the underlying patterns in the data.
  • Batch normalization plays a crucial role in stabilizing the training process. By standardizing the inputs to each layer, it helps mitigate the internal covariate shift problem, allowing for faster and more stable training [56].
  • The SiLU activation function introduces non-linearity into the network, which is essential for capturing complex relationships. SiLU has been shown to outperform ReLU in many scenarios [57]:
    $\mathrm{SiLU}(\mathbf{x})=\mathbf{x}\odot\sigma(\mathbf{x}),$ (8)
    where $\sigma(\mathbf{x})=(1+e^{-\mathbf{x}})^{-1}$ is the element-wise sigmoid function and $\odot$ is the element-wise product. The smooth nature of SiLU aids in the propagation of gradient during training.
  • Dropout serves as a powerful regularization step, preventing overfitting by randomly deactivating a fraction of nodes during training. This encourages the network to learn more robust features and reduces its reliance on any single node [58]. We set dropout rate of 0.5 for all Dropout layers in XAS-block.
  • The Softplus activation in the final layer ensures that the output respects the physical constraints of XAS spectra:
    $\mathrm{Softplus}(\mathbf{x})=\log(1+e^{\mathbf{x}}),$ (9)
    where all operations are applied element-wise. This function guarantees positive outputs, which is crucial as XAS intensities cannot be negative.
@@ II.5.1 Balanced Material Splitting
For training, we first divide the data into training, validation, and test sets, used respectively for model fitting, hyperparameter tuning, and final performance evaluation. The goal of our XAS prediction models is to be generalizable to unseen materials, not just unseen absorbing sites. To achieve this, we employ materials splitting, where all XAS spectra from a single material’s absorption sites are grouped together, ensuring they always fall into the same split. This ensures no data from any single material appears across multiple splits (e.g., a material with two atoms, in which one atom’s data is in the training set and the other atom in the testing set), preventing data leakage and overly optimistic performance estimates. We call this approach materials splitting, which provides a more stringent assessment of a model’s generalizability to new compounds.
@@ III.1 Spectra Visual Inspection
Figure 3: Heatmaps of FEFF XANES spectra. Color represents the density of spectral features, with red indicating a higher concentration of spectra and blue indicating a lower concentration (and white indicating no data).
@@
We proceed by direct visual inspection of the spectra database with focus first on the high-density region (dark red) of the entire FEFF database in Fig. 3. Ti and V XANES exhibit clearly visible pre-edge peaks. The pre-edge peak intensity decays when moving towards late 3d transition metals, consistent with previous observations [52]. Mn and Fe XANES show a much wider range in edge location between 10 to 20 eV than others, consistent with multiple commonly observed oxidation states. Mn XANES shows at least three edge positions as families of dark red curves and Fe XANES shows at least two with smaller spacing than Mn. Correspondingly, Mn and Fe spectra exhibit a broader spectral variation or data complexity than others. In contrast, Ni and Cu XANES show a relatively simple spectral shape with one dominant peak. Regarding Ti and Cu VASP database in Fig. 4, overall the spectral shape in VASP are similar to that in FEFF. However, the low density region (dark blue) in VASP has larger variations than FEFF in both elements.
@@ III.2 Transfer-Features Clustering
We select the node features of the absorbing site within the M3GNet latent space for transfer-feature extraction, as detailed in Sec. II.4.2. While this choice is intuitively appealing, it necessitates careful examination due to the feature smoothing phenomenon. Feature smoothing occurs when node features gradually become similar with each other in deeper layers of the model [63, 64]. This happens because the deeper convolutional layers increasingly expand the effective receptive field from which the features are being aggregated. In the case of M3GNet, feature-smoothing could potentially dilute local structural information when it passes through the M3GNet block, thereby hindering the site-specificity of predicted XAS and model performance.
@@
We plot two-dimensional UMAP of all transfer features in the FEFF ML dataset in Fig. 5a-d, colored by element type and local physical descriptors (OS, CN, and OCN). In doing so, we make two key observations as summarized below that provide compelling evidence that the latent-subspace-based transfer features can be used for XAS prediction without concern for potential feature smoothing inside the M3GNet block.
@@ III.3 Spectra Clustering
As shown in Fig. 5e, distinct clustering patterns form for each element. In addition, elements next to each other in the periodic table stay close together: early 3d transition metals (Ti, V, Cr and Mn) and late 3d transition metals (Fe, Co, Ni and Cu). However, there is not a clear separation between early and late 3d transition metals. The spectral similarity between different elements is supported by the observation in the raw spectra in Fig. 3, e.g., the pronounced pre-edge in Ti, V and Cr, and the single peak feature in Ni and Cu.
Further spectra clustering analysis was performed by labeling the UMAP pattern with physical descriptors (OS, CN and OCN) as shown in Fig. 5f-h. Within each element, one can see sub-level clustering with respect to the physical descriptors. For example, there is a dominant Ti4+ cluster (yellow) accompanied by a much smaller Ti2+ cluster (purple), while the Ti3+ cluster (red) is less significant and more scattered. This is consistent with the fact that the Ti database is dominated by Ti4+ cations. Furthermore, there is a clear distinction in Mn and Fe clusters with respect to OS (2+, 3+ and 4+), resonating the strong trend in edge positions in the raw spectra in Fig. 3, because edge position is a well established proxy of ion OS. Similarly, sub-regions of CN can be seen within each cluster, e.g. 4-, 5- and 6-coordinated Ti and V with 6-coordinated motifs being most prevalent, consistent with previous work based on PCA analysis [52]. OCN is a more subtle physical descriptor, which is correlated to main peak intensity [53]. However, such trend is not obvious in the raw spectra, as it is entangled with spectral variations caused by other physical descriptors [53]. It is quite remarkable that UMAP can in fact form distinct clusters of OCN, although its pattern is more complex than OS and CN.
@@ IV.3 Performance Overview
Table 1: Performance metric of ExpertXAS ($\eta_{\text{E}}$), UniversalXAS ($\eta_{\text{U}}$), and Tuned-UniversalXAS ($\eta_{\text{T}}$) models. The best performance for each element is highlighted in bold. The improvement column shows the percentage increase in performance of the Tuned-UniversalXAS model over the ExpertXAS model. The Win Rate shows improvement across across spectra defined by Eq. 17 and energy grid defined by Eq. 18.
@@ IV.3.2 UniversalXAS Performance
The MSE distribution of ExpertXAS models shown in Fig. 8 demonstrates consistent patterns across different elements when examined on a logarithmic scale (log10(MSE)). These errors manifest as bell-shaped distributions centered between -9 and -7, where smaller errors are common and larger MSE become exponentially rare. While spanning approximately three orders of magnitude (-10 to -7), these distributions consistently show lower errors than their respective baselines. The similarity in distribution shapes and their consistent positioning relative to the baselines across the transition metal series suggests robust model behavior.
@@ IV.4 Example Predictions
Figure 9: Spectra from deciles $D_{i}$ sorted by the MSE-per-spectra of the best model’s (Tuned-Universal) prediction. Solid lines represent simulated spectra, dashed lines represent predicted spectra (ground truth), and shaded areas highlight the differences. Performance $\eta$ for Mn, Fe and Cu corresponds to the maximum, median, and minimum $\eta_{i}$ across elements.
We now provide specific examples to demonstrate the predictive power of XASModels. Fig. 9 shows spectrum predictions for Mn, Fe, and Cu, corresponding to the highest, median, and lowest performing models, respectively, within the Tuned-UniversalXAS variant. Predictions are categorized into deciles based on MSE per spectrum, with $D_{i}$ representing the $i$-th percentile of the MSE distribution. We also provide comprehensive view of example predictions from nine deciles for each of the ten element/simulation pairs, totaling 252 spectra in Supplementary Figs. S1-S3.
As shown in Fig. 9, in the best performing XASModel of Mn, the predicted spectra closely align with the simulated spectra across all deciles. The shaded areas, representing differences between predicted and simulated spectra, are minimal even in the lower deciles. For the representative median performance of Fe, predictions show strong agreement in the higher deciles ($D_{10}$-$D_{50}$), with slightly larger deviations in the lower deciles, particularly in the main-edge and post-edge regions. For the case of Cu, which represents the model with lowest performance, predictions still accurately capture the main edge and overall spectral shape in the higher deciles, with more noticeable differences in the fine details of the lower deciles.
@@ IV.5.1 Improvement Distribution
$\Delta_{j}^{(i)}\equiv|\hat{\mu}_{j}^{(i,\text{expert})}-\mu_{j}^{(i)}|-|\hat{\mu}_{j}^{(i,\text{tuned})}-\mu_{j}^{(i)}|.$ (15)
This metric provides a point-by-point comparison of model performance. Cases where $\Delta_{j}^{(i)}>0$ indicate superior performance by the Tuned-UniversalXAS model, while $\Delta_{j}^{(i)}<0$ favors the ExpertXAS model. Fig. 10 illustrates the distribution of all residual differences partitioned by its sign. We observe dominant instances of performance improvements (darker bars) compared to performance degradations (lighter bars) for most elements. Notably, for elements like Mn and Cu, where the difference in $\eta$ values is small, we see a more balanced distribution between improvements and degradations. These observations aligns with earlier observations in Sec. IV.3 using the $\eta$ metric.
Figure 10: Distribution of residual differences $\Delta=\{\Delta_{j}^{(i)}\mid i=1,2,\ldots,N;\,j=1,2,\ldots,M\}$ between Tuned-UniversalXAS and ExpertXAS predictions (see Eq. 15). For easy comparison between improvements and degradations, the data is partitioned by sign of $\Delta_{j}^{(i)}$ and plotted against its absolute value in logarithmic scale. Darker bars ($\Delta_{j}^{(i)}>0$) indicate predictions where Tuned-UniversalXAS outperformed ExpertXAS, while lighter bars ($\Delta_{j}^{(i)}<0$) indicate the opposite. Bar heights represent the percentage of the data points, with the total number of data points given by the product of the number of energy grids and the spectra count.

### v3

# OmniXAS: A Universal Deep-Learning Framework for Materials X-ray Absorption Spectra

@@ Abstract
X-ray absorption spectroscopy (XAS) is a powerful characterization technique for probing the local chemical environment of absorbing atoms. However, analyzing XAS data presents with significant challenges, often requiring extensive, computationally intensive simulations, as well as significant domain expertise. These limitations hinder the development of fast, robust XAS analysis pipelines that are essential in high-throughput studies and for autonomous experimentation. We address these challenges with OmniXAS, a framework that contains a suite of transfer learning approaches for XAS prediction, each uniquely contributing to improved accuracy and efficiency, as demonstrated on K-edge spectra database covering eight 3d transition metals (Ti – Cu). The OmniXAS framework is built upon three distinct strategies. First, we use M3GNet [Nat. Comput. Sci., 2, 718 (2022)] to derive latent representations of the local chemical environment of absorption sites as input for XAS prediction, achieving significant improvements over conventional featurization techniques. Second, we employ a hierarchical transfer learning strategy, training a universal multi-task model across elements before fine-tuning for element-specific predictions. Models based on this cascaded approach after element-wise fine-tuning outperform element-specific models by up to 69%. Third, we implement cross-fidelity transfer learning, adapting a universal model to predict spectra generated by simulation of a different fidelity with a much higher computational cost. This approach improves prediction accuracy by up to 11% over models trained on the target fidelity alone. Our approach significantly boosts the throughput of XAS modeling by orders of magnitude as compared to first-principles simulations and is extendable to XAS prediction for a broader range of elements. The proposed transfer learning framework is generalizable to enhance deep-learning models that target other properties in materials research.
@@ I Introduction
A fast and accurate spectral prediction surrogate model can play an important role in the XAS analysis pipeline. For a given target measured XAS spectrum, one can use an efficient search algorithm to sample the relevant material space and identify candidate structures. The corresponding XAS spectra are compared with experiment to identify the best match. For complex material systems (e.g., amorphous materials, defects, nanostructures, liquid systems and interfaces), the material space can be very large and each candidate structure may contain hundreds or thousands of atoms. It becomes impractical to perform a huge number of first-principles XAS simulations on very large systems in a high-throughput manner. A high-fidelity XAS surrogate model is the solution to this problem. The significance of this approach was demonstrated in the recent study of amorphous carbon [25], where a diffusion model for structure sampling was combined with an XAS surrogate model to analyze the measured C K-edge XANES spectra.
In this work, we address the need for a generalizable ML framework to predict XANES spectra from atomic structures. Our approach leverages transfer learning and domain adaptation techniques to develop universal models. By adopting M3GNet-derived latent features [28] to predict K-edge XANES spectra for a range of 3d transition metals, we demonstrate that this method not only outperforms traditional featurization techniques but also effectively bridges different element types, providing a scalable solution for XANES analysis across a wide range of materials. By first training a universal model at a fidelity with low computational cost, we demonstrate that ML models at computationally expensive fidelities can be efficiently developed by fine-tuning the universal model. The OmniXAS framework lays the groundwork for future applications in high-throughput and real-time XAS analysis, significantly reducing the reliance on computationally expensive simulations.
@@ II.1 Data Acquisition and Curation
Before ML models can be trained, structure and spectral data need to be sourced, curated, and standardized. We used the Lightshow Python package [13, 10] to pull all available Materials Project structural data of pure metals and primary, secondary, and ternary oxides via Pymatgen [31] for Ti, V, Cr, Mn, Fe, Co, Ni and Cu. FEFF9 and VASP spectral input files were also written using Lightshow.
@@ II.4.1 M3GNet block
M3GNet is a graph deep learning model designed to predict the inter-atomic potential, forces, and stresses from the material graph representation [28]. It was trained on a comprehensive dataset from the Materials Project, comprising over 17 million entries and encompassing 89 elements from the periodic table. The efficiency of M3GNet was demonstrated by the study that predicted almost 2 million stable materials from a total of approximately $32$ million candidate structures [28].
@@ II.4.3 XAS block
Here, $\hat{\bm{\mu}}\in\mathbb{R}^{141}$ represents the predicted XAS spectrum across 141 energy grid points. Each component of the MLP serves a specific purpose in the learning process, contributing to the model’s overall effectiveness. Further details of the MLP can be found in the Supplementary Materials.
@@ II.5.1 Balanced Material Splitting
For training, we first divide the data into training, validation, and test sets, used respectively for model fitting, hyperparameter tuning, and final performance evaluation. The goal of our XAS prediction models is to be generalizable to unseen materials, not just unseen absorbing sites. To achieve this, we employ materials splitting, where all XAS spectra from a single material’s absorption sites are grouped together, ensuring they always fall into the same split. This ensures no data from any single material appears across multiple splits (e.g., a material with two unique absorbing sites, in which one site’s data is in the training set and the other site in the testing set), preventing data leakage and overly optimistic performance estimates. We call this approach materials splitting, which provides a more stringent assessment of a model’s generalizability to new compounds.
@@ III.1 Spectra Visual Inspection
Figure 3: Heatmaps of FEFF XANES spectra. Color represents the density of spectral features.
@@
We proceed by direct visual inspection of the spectra database with focus first on the high-density region (dark red) of the entire FEFF database in Fig. 3. Ti and V XANES exhibit clearly visible pre-edge peaks. The pre-edge peak intensity decays when moving towards late 3d transition metals, consistent with previous observations [52]. Mn and Fe XANES show a much wider range in edge location than others, consistent with multiple commonly observed oxidation states. Mn XANES shows at least three edge positions as families of dark red curves and Fe XANES shows at least two with smaller spacing than Mn. Correspondingly, Mn and Fe spectra exhibit a broader spectral variation or data complexity than others. In contrast, Ni and Cu XANES show a relatively simple spectral shape with one dominant peak. Regarding Ti and Cu VASP database in Fig. 4, overall the spectral shape in VASP are similar to that in FEFF. However, the low density region (dark blue) in VASP has larger variations than FEFF in both elements.
@@ III.2 Transfer-Features Clustering
We select the node features of the absorbing site within the M3GNet latent space for transfer-feature extraction, as detailed in Sec. II.4.2. While this choice is intuitively appealing, it necessitates careful examination due to the feature smoothing phenomenon. Feature smoothing occurs when node features gradually become similar with each other in deeper layers of the model [60, 61]. This happens because the deeper convolutional layers increasingly expand the effective receptive field from which the features are being aggregated. In the case of M3GNet, feature-smoothing could potentially dilute local structural information when it passes through the M3GNet block, thereby hindering the site-specificity of the predicted XAS and model performance.
@@
We plot two-dimensional UMAP of all transfer features in the FEFF ML dataset in Fig. 5a-d, colored by element type and local physical descriptors (OS, CN, and OCN). In doing so, we make two key observations as summarized below that provide compelling evidence that the latent-subspace-based transfer features can be used for XAS prediction without concern for excessive feature smoothing inside the M3GNet block.
@@ III.3 Spectra Clustering
As shown in Fig. 5e, distinct clustering patterns form for each element. In addition, elements next to each other in the periodic table stay close together: early 3d transition metals (Ti, V, Cr and Mn) and late 3d transition metals (Fe, Co, Ni and Cu). However, there is not a clear separation between early and late 3d transition metals. The spectral similarity between different elements is supported by the observation in the raw spectra in Fig. 3, e.g., the pronounced pre-edge in Ti, V and Cr, and the single peak feature in Ni and Cu. The distinction between different elements observed in the spectral UMAP may arise from several factors, including differences in core-hole lifetimes and absorber atom phase shifts, as well as variations in scattering path lengths caused by differing bond lengths under the same local symmetry.
Further spectra clustering analysis was performed by labeling the UMAP pattern with physical descriptors (OS, CN and OCN) as shown in Fig. 5f-h. Within each element, one can see sub-level clustering with respect to the physical descriptors. For example, there is a dominant Ti4+ cluster (yellow) accompanied by a much smaller Ti2+ cluster (purple), while the Ti3+ cluster (red) is less significant and more scattered. This is consistent with the fact that the Ti database is dominated by Ti4+ cations. Furthermore, there is a clear distinction in Mn and Fe clusters with respect to OS (2+, 3+ and 4+), reflecting the clear trend in edge positions in the raw spectra in Fig. 3, because edge position is a well established proxy of ion OS. Similarly, sub-regions of CN can be seen within each cluster, e.g. 4-, 5- and 6-coordinated Ti and V with 6-coordinated motifs being most prevalent, consistent with previous work based on PCA analysis [52]. OCN is a more subtle physical descriptor, which is correlated to main peak intensity [53]. However, such trend is not obvious in the raw spectra, as it is entangled with spectral variations caused by other physical descriptors [53]. It is quite remarkable that UMAP can in fact form distinct clusters of OCN, although its pattern is more complex than OS and CN.
@@ IV.3 Performance Overview
Table 1: Performance metric of ExpertXAS ($\eta_{\text{E}}$), UniversalXAS ($\eta_{\text{U}}$), and Tuned-UniversalXAS ($\eta_{\text{T}}$) models. The best performance for each element is highlighted in bold. The improvement column shows the percentage increase in performance of the Tuned-UniversalXAS model over the ExpertXAS model. The Win Rate shows improvement across spectra defined by Eq. 14 and energy grid defined by Eq. 15.
@@ IV.3.1 ExpertXAS Performance
The MSE distribution of ExpertXAS models shown in Fig. 8 demonstrates consistent patterns across different elements when examined on a logarithmic scale (log10(MSE)). These errors exhibit bell-shaped distributions centered between -9 and -7, except for Mn FEFF, which shows a bimodal distribution. Small errors are common, while large errors become exponentially rare. While spanning approximately three orders of magnitude, these distributions consistently show lower errors than their respective baselines. The similarity in distribution shapes and their consistent positioning relative to the baselines across the transition metal series suggests robust model behavior.
@@ IV.3.3 Tuned-UniversalXAS Performance
Figure 9: Spectra from deciles $D_{i}$ sorted by the MSE-per-spectra of the best model’s (Tuned-Universal) prediction. Solid lines represent simulated spectra (ground truth), dashed lines represent predicted spectra, and shaded areas highlight the differences. Performance of Mn, V and Cu corresponds to the maximum, median, and minimum $\eta_{i}$ across elements.
@@ IV.4 Example Predictions
We now provide specific examples to demonstrate the predictive power of XASModels. Fig. 9 shows FEFF spectral predictions for Mn, V, and Cu, corresponding to the highest, median, and lowest performing models, respectively, within the Tuned-UniversalXAS variant. Predictions are categorized into deciles based on MSE per spectrum, with $D_{i}$ representing the $i$-th percentile of the MSE distribution. We also provide comprehensive view of example predictions from nine deciles for each of the ten element/simulation pairs, totaling 252 spectra in Supplementary Figs. S1-S3.
[…]
@@
Given the ultimate goal of ML models as an accurate and efficient tool for analyzing experimental XANES spectra, it is important to understand how accurately ML predictions can reproduce experiment. However, we note that the OmniXAS framework developed here is designed to serve as a surrogate for XANES simulations, and as such, the models are not trained to accurately reproduce experimental data to any higher accuracy than FEFF or VASP. Since the quantitative assessment of OmniXAS against the ground truth of the simulations has been rigorously established, a systematic benchmark of simulation (e.g., FEFF and VASP) against experiment is warranted for future studies.
@@ IV.5.1 Improvement Distribution
$\Delta_{j}^{(i)}\equiv|\hat{\mu}_{j}^{(i,\text{E})}-\mu_{j}^{(i)}|-|\hat{\mu}_{j}^{(i,\text{T})}-\mu_{j}^{(i)}|.$ (12)
This metric provides a point-by-point comparison of model performance. Cases where $\Delta_{j}^{(i)}>0$ indicate superior performance by the Tuned-UniversalXAS model, while $\Delta_{j}^{(i)}<0$ favors the ExpertXAS model. Fig. 10 illustrates the distribution of all residual differences partitioned by its sign. We observe dominant instances of performance improvements (darker bars) compared to performance degradations (lighter bars) for most elements. Notably, for Cu, where the difference in $\eta$ value is small, we see a more balanced distribution between improvements and degradations. These observations aligns with earlier observations in Sec. IV.3 using the $\eta$ metric.
Figure 10: Distribution of residual differences $\Delta=\{\Delta_{j}^{(i)}\mid i=1,2,\ldots,N;\,j=1,2,\ldots,M\}$ between Tuned-UniversalXAS and ExpertXAS predictions (see Eq. 12). For easy comparison between improvements and degradations, the data is partitioned by sign of $\Delta_{j}^{(i)}$ and plotted against its absolute value in logarithmic scale. Darker bars ($\Delta_{j}^{(i)}>0$) indicate predictions where Tuned-UniversalXAS outperforms ExpertXAS, while lighter bars ($\Delta_{j}^{(i)}<0$) indicate the opposite. Bar heights represent the percentage of the data points, with the total number of data points given by the product of the number of energy grids and the spectra count.